# Deep Reinforcement Learning for Controlled Traversing of the Attractor Landscape of Boolean Models in the Context of Cellular Reprogramming

### v2

@@ Abstract
Cellular reprogramming can be used for both the prevention and cure of different diseases. However, the efficiency of discovering reprogramming strategies with classical wet-lab experiments is hindered by lengthy time commitments and high costs. In this study, we develop a novel computational framework based on deep reinforcement learning that facilitates the identification of reprogramming strategies. For this aim, we formulate a control problem in the context of cellular reprogramming for the frameworks of BNs and PBNs under the asynchronous update mode. Furthermore, we introduce the notion of a pseudo-attractor and a procedure for identification of pseudo-attractor state during training. Finally, we devise a computational framework for solving the control problem, which we test on a number of different models.
@@ 1 Introduction
Complex diseases pose a great challenge largely because genes and gene products operate within a complex system – the gene regulatory network (GRN). There is an inherent dynamic behaviour emerging from the structural wiring of a GRN: gene expression profiles, i.e., states of a GRN, evolve in time to finally reach stable states referred to as attractors. Attractors correspond to cell types or cell fates [13]. During normal development of a multi-cellular organism, not all attractors are manifested. Some of the ‘abnormal attractors’, associated with diseases, become accessible by disturbance of the GRN’s dynamics. This is seldom a consequence of a disruption in a single gene, but rather arises as an aftermath of GRN perturbations [3]. This could be cured by guiding cells to desired ‘healthy’ attractors with experimental techniques of cellular reprogramming, i.e., the artificial changing of cell fate. Unfortunately, finding effective interventions that trigger desired changes using solely wet-lab experiments is difficult, costly, and requires lengthy time commitments. This motivates us to consider in-silico approaches.
Although various computational frameworks are commonly used to model GRNs, the formalism of Boolean networks (BNs) and its extension, i.e., probabilistic Boolean networks (PBNs), have the advantage of being simple yet capable of capturing the important dynamic properties of the system under study. As such, they facilitate the modelling of large biological systems. This is especially relevant in the context of in-silico identification of effective cellular reprogramming strategies, which requires large GRNs to be modelled.
Identification of cellular reprogramming strategies can be stated as a control problem of BN and PBN models of GRNs. Although many BN/PBN control methods exist in the literature, the existing structure- and dynamics-based state-of-the-art computational techniques are limited to small and mid-size networks, i.e., of up to a hundred of genes or so, usually requiring the systems to be decomposed in some way. This is often insufficient for cellular reprogramming considerations.
The issue of scalability can be addressed by devising new methods based on deep reinforcement learning (DRL) techniques, which have proved very successful in decision problems characterised by huge state-action spaces. To contribute to the realisation of this idea, we formulate a control problem in the context of cellular reprogramming for the frameworks of BNs and PBNs under the asynchronous update mode. Furthermore, we introduce the notion of a pseudo-attractor and a procedure for identifying pseudo-attractor states during DRL agent training. Finally, these contributions allow us to devise a DRL-based framework for solving the control problem. We consider our contributions as a relevant step towards achieving scalable control methods for large Boolean models of GRNs for identifying effective and efficient cellular reprogramming strategies.
The paper is structured as follows. Related work is discussed in Sec. 2. Preliminaries are provided in Sec. 3. We formulate our control problem in the context of cellular reprogramming in Sec. 4 and devise our DRL-based control framework in Sec. 5. The experiments performed to evaluate our framework and the obtained results are presented in Sec.6 and Sec. 7, respectively. Finally, we conclude our study in Sec. 8.
2 Related work
2.1 Dynamics-based approaches to GRN control
@@
Identification of proper control strategies for non-linear systems requires both network structure and dynamics [11]. Thus, we focus on dynamic-based and DRL-based methods for BN/PBN control. An efficient method based on the ‘divide and conquer’ strategy was proposed in [18] to solve the minimal one-step source-target control problem by using instantaneous, temporary, and permanent gene perturbations. The minimal sequential source-target control and the target control problems of BNs were considered in [23] and [22], respectively. All these methods were implemented as a software tool CABEAN [24]. Recently, semi-symbolic algorithms were proposed in [4] to stabilise partially specified asynchronous BNs in states exhibiting specific traits. In [19], the control problem for the most permissive BN update mode in the context of fixed points and minimal trap spaces is considered.
2.2 DRL-based approches to GRN control
@@
The application of reinforcement learning for controlling GRNs was pioneered by in [21] with focus on how to control GRNs by avoiding undesirable states in terms of steady state probabilities of PBNs. The main idea was to treat the time series gene expression samples as a sequence of experience tuples and use a batch version of Q-Learning to produce an approximated policy over the experience tuples. Later, the BOAFQI-Sarsa method that does not require time series samples was devised in [16]. A batch reinforcement learning method, mSFQI, was proposed in [15] for control based on probabilities of gene activity profiles. Recently, the study of [1] used a Deep Q-Network with prioritised experience replay, for control of synchronous PBNs to drive the networks from a particular state towards a more desirable one. Finally, a DRL-based approximate solution to the control problem in synchronous PBNs was proposed in [14]. The proposed method finds a control strategy from any network state to a specified target attractor using a Double Deep Q-Network model.
3 Preliminaries
3.1 Boolean and probabilistic Boolean networks
Boolean networks is a well established framework for the modelling of GRNs. A Boolean network consists of nodes, that can be in one of two states, and functions describing how the individual nodes interact with each other. PBNs are an extension of the formalism of BNs.
Definition 3.1.
(Boolean Network) A Boolean Network is defined as a pair $(V,F)$, where $V=\{x_{1},x_{2},\ldots,x_{n}\}$ is a set of binary-valued nodes (also referred to as genes) and $F=\{f_{1},f_{2},\ldots,f_{n}\}$ is a set of Boolean predictor functions, where $f_{i}(x_{i_{1}},x_{i_{2}},\ldots,x_{i_{k}})$ defines the value of node $x_{i}$ depending on the values of the $k\leq n$ parent nodes $x_{i_{1}},x_{i_{2}},\ldots,x_{i_{k}}$ with $i_{j}\in[1..n]$ for $j\leq k$.
Since interactions in biology are usually more complex we need a more general model of a GRN. We achieve it by allowing for each node to have multiple Boolean functions. Formally, probabilistic Boolean networks are defined as follows:
Definition 3.2.
[…]
A PBN in which each node only admits only one Boolean function is a Boolean network.
3.2 Network dynamics
@@
We define a state of a BN/PBN as an $n$-dimensional vector $\mathbf{s}\in\{0,1\}^{n}$, where the $i$-th element represents the state of gene $x_{i}$ for $i\in[1..n]$. A BN/PBN evolves in discrete time steps. It starts in an initial state $\mathbf{s}_{0}$ and its state gets updated in every time step in accordance to the predictor functions. In this study, we focus on the asynchronous updating, which is preferable in the context of GRN modelling. Under the asynchronous scheme, a single gene $x_{i}$ is selected and updated in accordance with its predictor function $f_{i}$ (BNs) or one randomly selected from $F_{i}$ in accordance with $\mathbf{c}^{i}$. The network dynamics can be depicted in the form of a state transition graph. Based on this concept of, we can introduce the notion of a BN/PBN attractor.
Definition 3.3.

### v3

# pbn-STAC: Deep Reinforcement Learning-based Framework for Cellular Reprogramming

@@ Abstract
Cellular reprogramming can be used for both the prevention and cure of complex diseases. However, the efficiency of discovering reprogramming strategies with classical wet-lab experiments is hindered by lengthy time commitments and high costs. In this study, we leverage deep reinforcement learning to develop a novel computational framework that facilitates the identification of reprogramming strategies. To this aim, we formulate a control problem in the context of cellular reprogramming for the Boolean and probabilistic Boolean network models of gene regulatory networks under the asynchronous update mode. Furthermore, to facilitate scalability, we introduce the notion of a pseudo-attractor and a procedure for the identification of pseudo-attractor states. Finally, we devise a computational framework for solving the control problem, which we test on a number of biological networks.
@@ 1 Introduction
Complex diseases pose a great challenge largely because genes and gene products operate within a complex system – the gene regulatory network (GRN). There is an inherent dynamic behaviour emerging from the structural wiring of a GRN: gene expression profiles, i.e., states of a GRN, evolve in time to finally reach stable states referred to as attractors. Attractors correspond to cell types or cell fates [12]. During normal development of a multi-cellular organism, not all attractors are manifested. Some of the ‘abnormal attractors’, associated with diseases, become accessible by disturbance of the GRN’s dynamics. This is seldom a consequence of a disruption in a single gene, but rather arises as an aftermath of GRN perturbations [3]. This could be cured by guiding cells to desired ‘healthy’ attractors with experimental procedures of cellular reprogramming. Unfortunately, finding effective interventions that trigger desired changes using solely wet-lab experiments is difficult, costly, and requires lengthy time commitments. This motivates us to consider in-silico approaches. In the context of computational modelling of GRNs, cellular reprogramming can be stated as a control problem. Since GRNs are complex systems characterised by emergence, computing cellular reprogramming procedures requires large GRN models to be considered. The modelling frameworks of Boolean networks (BNs) and probabilistic Boolean networks (PBNs) are well-suited for the holistic modelling of GRNs due to the advantage of being simplistic generalisations of the system’s dynamics and yet capable of capturing the intrinsic non-linear dynamic properties.
@@
Computation of proper control strategies for non-linear systems requires both network structure and dynamics [9]. Recently, a number of state-of-the-art methods have been proposed. An efficient method based on the ‘divide and conquer’ strategy was proposed in [18] to solve the minimal one-step source-target control problem by using instantaneous, temporary, and permanent gene perturbations. The minimal sequential source-target control and the target control problems of BNs were considered in [23] and [22], respectively. All these methods were implemented in the CABEAN [24] tool. Recently, semi-symbolic algorithms were proposed in [4] to stabilise partially specified asynchronous BNs in states exhibiting specific traits. In [19], the control problem for the most permissive BN update mode in the context of fixed points and minimal trap spaces was considered.
The existing structure- and dynamics-based state-of-the-art computational techniques are limited to small and mid-size networks due to the infamous state space explosion problem. They usually require the systems to facilitate some kind of decomposition. This however is often too restrictive for cellular reprogramming considerations.
@@
The issue of scalability can be addressed by devising new methods based on deep reinforcement learning (DRL) techniques, which have proved very successful in decision problems characterised by huge state-action spaces. The application of reinforcement learning for controlling GRNs was pioneered in [21] with focus on how to control GRNs by avoiding undesirable states in terms of steady state probabilities of PBNs. The main idea was to treat the time series gene expression samples as a sequence of experience tuples and use a batch version of Q-Learning to produce an approximated policy over the experience tuples. Later, the BOAFQI-Sarsa method that does not require time series samples was devised in [17]. A batch reinforcement learning method, mSFQI, was proposed in [16] for control based on probabilities of gene activity profiles. Recently, the study of [1] used a Deep Q-Network with prioritised experience replay, for control of synchronous PBNs to drive the networks from a particular state towards a more desirable one. Finally, a DRL-based approximate solution to the control problem in synchronous PBNs was proposed in [15]. The proposed method finds a control strategy from any network state to a specified target attractor using a Double Deep Q-Network.
This study represents an exploration into the potential of DRL for developing scalable control methods for both BN and PBN models of GRNs in the context of cellular reprogramming. Our contributions are as follows. We formulate a novel control problem, i.e., the source-target attractor control for BN/PBN models of GRNs, tailored for cellular reprogramming in the sense that control interventions are allowed to take place only in states corresponding to phenotypic states of the cell. Moreover, in contrast to the existing studies on DRL-based control methods, where synchronous update scheme was considered, our focus is on the computationally more challenging asynchronous mode since it is commonly considered more appropriate for the modelling of GRNs. Next, as identifying attractors in large networks is ineffective by itself, we introduce the notion of a pseudo-attractor and develop a procedure for identifying pseudo-attractor states during DRL agent training. We then devise pbn-STAC, i.e., a DRL-based framework for solving the source-target attractor control problem, which we test on a number of case studies. In order for the predictions on the control strategies to be more realistic in terms of wet-lab practices, the control actions are restricted to be applied in (pseudo-)attractor states only. We consider our contributions as a relevant step towards development of scalable computational methods based on DRL for the control of BN/PBN models of GRNs in the context of cellular reprogramming. We present the materials, methods, and experimental results of pbn-STAC in the subsequent sections.
2 Materials and methods
2.1 Boolean networks
A PBN is a discrete dynamical system defined as a pair $(V,\mathcal{F})$, where $V=\{x_{1},x_{2},\ldots,x_{n}\}$ is a set of binary-valued nodes (also referred to as genes) and $\mathcal{F}=(F_{1},F_{2},\ldots,F_{n})$ is a list of sets. Each node $x_{i}\in V$ has associated a set $F_{i}\in\mathcal{F}$ of predictor functions, i.e., $F_{i}=\{f^{i}_{1},f^{i}_{2},\ldots,f^{i}_{l(i)}\}$, where $l(i)$ is the number of predictor functions of node $x_{i}$. Each $f^{i}_{j}\in F_{i}$ is a Boolean function defined with respect to a subset of $V$ denoted $\textrm{Pa}(f^{i}_{j})$ and referred to as parent nodes of $f^{i}_{j}$. For each $x_{i}\in V$, there is a probability distribution $\mathbf{c}^{i}=(c^{i}_{1},c^{i}_{2},\ldots,c^{i}_{l(i)})$ on $F_{i}$, where each predictor function $f^{i}_{j}\in F_{i}$ has an associated selection probability denoted $c^{i}_{j}$; it holds that $\sum_{j=1}^{l(i)}c^{i}_{j}=1$. A BN is a PBN in which each node admits only a single predictor function.
@@ 2.1 Boolean networks
We define a state of a BN/PBN as an $n$-dimensional vector $\mathbf{s}\in\{0,1\}^{n}$, where the $i$-th element represents the state of gene $x_{i}\in V$. A BN/PBN evolves by starting in an initial state and updating it at discrete time steps in accordance with the selected predictor functions. Herein, we focus on the asynchronous updating scheme, where at each time step the state value for a single gene $x_{i}$ is updated in accordance with a randomly selected predictor function from $F_{i}$ in accordance with $\mathbf{c}^{i}$.